there are 3 liquids a,b,c of equal masses, they are at temprature 10,15,25centigrade when liquid a and b are mixed ther common temprature is 12centigrade and b and c are mixed then it is 21centigrade. what will be the common temprature of a and c.


 

Grade:

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear saurabh

let  c1 ,c2,c3 are the heat capacity of  liquid a,b,c

when a and b are mixed.

mc1*10  + mc2*15 =mc1* 12 + mc2*12

 3c2 =  2c1.............1

 

when b and c are mixed.

mc2*15  + mc3*25 =mc2* 21 + mc3*21

 2c3 =  3c2.............2

from equation 1 and 2

c1=c3

now liquid a and c are mixed  let the final temperature is T

mc1*10 + mc3*25 = mc1*T + mc3*T

 T = 17.5

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin

Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

let c1 ,c2,c3 are the heat capacity of liquid a,b,c

when a and b are mixed.

mc1*10 + mc2*15 =mc1* 12 + mc2*12

3c2 = 2c1.............1



when b and c are mixed.

mc2*15 + mc3*25 =mc2* 21 + mc3*21

2c3 = 3c2.............2

from equation 1 and 2

c1=c3

now liquid a and c are mixed let the final temperature is T

mc1*10 + mc3*25 = mc1*T + mc3*T

T = 17.5

Thanks and Regards